OpenStudy (anonymous):

Please help! Multiply and simplify: 5x+15 x^3 ------ * ---- x x+3

OpenStudy (anonymous):

I just wanted to reply to say that your username is cool. Multiply each side by the opposite numerator

OpenStudy (vishweshshrimali5):

Hey its easy dude

OpenStudy (anonymous):

@ilovemath Haha :) Thank you. You too :P Thanks for the help!

OpenStudy (vishweshshrimali5):

See, (5x + 15) (x^3) --------------- x(x+3) = 5(x+3)(x^3) ----------- x (x+3) = 5x^(3-1) = 5x^2 This is your answer

OpenStudy (vishweshshrimali5):

Are you satisfied with the answer

OpenStudy (vishweshshrimali5):

Hey thanx for giving me a medal friend

OpenStudy (anonymous):

Thank YOU for the help :)

OpenStudy (vishweshshrimali5):

Mention not friend